OpenStudy (anonymous):

Solve the combined inequality and select the answer that best describes the graph of the solution. 4x + 6 < 14 or 8x - 6 > 26

OpenStudy (kamille):

4x + 6 < 14 4x<14-6 4x<8|:4 x<2 \[(-\infty;2)\]

OpenStudy (anonymous):

Solve inequalities just like any other equation. 4x + 6 < 14 4x < 8 x < 2 8x - 6 > 26 8x > 32 x > 4 \[x \in ( - \infty ,2) \cup (4,\infty )\]